OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10 Answer x = 8 x = 2 x = -8 x = -2

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

x-3= 10/-2 x=-5 +3 x=-2